(b) What is the acceleration of gravity on the surface (or outer limit) of Neptune? The mass of Neptune is 1.02 1026 kg and its radius is 2.48 107 m.

m/s^2
Answer #1

To find the magnitude of the acceleration of gravity on the surface of a body with a big mass against any mass on it's surface the acceleration can be written as follows:


g= \frac{G\cdot M}{r^{2}}

We must also know that gravity tends to attract bodies so the gravitational force point to the center of the planet, in this case, Neptune.

Substituting values we have:


g= \frac{6.67384\cdot 10^{-11}Nm^{2}/kg^{2}\cdot 1.02\cdot 10^{26}kg}{(2.48\cdot 10^{7})^{2}}=11.06 m/s^{2}

This acceleration is bigger than the 9.81 m/s2 on earth. This also means that your weight would be higher on Neptune than on earth since this force is proportional to the acceleration of gravity